# Template Free Reconstruction of Human-object Interaction with Procedural Interaction Generation

### 12 Dec 2023 v1

@@ Abstract
Reconstructing human-object interaction in 3D from a single RGB image is a challenging task and existing data driven methods do not generalize beyond the objects present in the carefully curated 3D interaction datasets. Capturing large-scale real data to learn strong interaction and 3D shape priors is very expensive due to the combinatorial nature of human-object interactions. In this paper, we propose ProciGen (Procedural interaction Generation), a method to procedurally generate datasets with both, plausible interaction and diverse object variation. We generate 1M+ human-object interaction pairs in 3D and leverage this large-scale data to train our HDM (Hierarchical Diffusion Model), a novel method to reconstruct interacting human and unseen object instances, without any templates. Our HDM is an image-conditioned diffusion model that learns both realistic interaction and highly accurate human and object shapes. Experiments show that our HDM trained with ProciGen significantly outperforms prior methods that require template meshes, and our dataset allows training methods with strong generalization ability to unseen object instances. Our code and data will be released at: https://virtualhumans.mpi-inf.mpg.de/procigen-hdm.
@@ 1 Introduction
Modelling interactions between humans and their surroundings is important for applications like creating realistic avatars, robotic control and gaming. In this paper, we address the task of jointly reconstructing human and object from a monocular RGB image, without any prior object templates. This is very challenging due to depth-scale ambiguity, occlusions, diverse human pose and object shape variations. Data-driven methods have shown great progress in reconstructing humans [65, 66, 67, 42, 38] or objects [51, 100] from monocular inputs thanks to large-scale datasets [9, 35, 59, 1, 96, 19, 12, 86]. However, methods for joint interaction reconstruction are still constrained by the amount of available data. Recent datasets like BEHAVE [7], InterCap [33] capture real interactions with 10 to 20 different objects, which is far away from the number of objects in reality: the chair category from ShapeNet [12] alone has more than 6k different shapes. Training on these real datasets has limited generalization ability to unseen objects (Sec. 4.3). Capturing real interaction data with more objects is prohibitively expensive due to the combinatorial nature: the number of humans times the number of objects leads to a huge number of variations. This motivates us to generate synthetic data which has been shown effective for pre-training reconstruction methods [9, 59, 28, 66, 51].
@@
Synthesizing realistic interaction for different objects is non-trivial due to variations of object topology, geometry details and complex interaction patterns. To address this, we propose Procedural interaction Generation (ProciGen), a method to generate interaction data with diverse object shapes. We design our method based on the key idea that the way humans interact with objects of the same category is similar. And despite the geometry variations, one can still establish semantically meaningful correspondence between different objects. More specifically, we train an autoencoder to obtain correspondences between different objects from the same category, which are then used to transfer contacts from already captured human-object interactions to new object instances. Our method is scalable and allows the multiplicative combination of datasets to generate over a million interactions with more than 21k different object instances, which is not possible via real data capture.
@@
Current reconstruction methods [87, 88, 7, 56] are not only bottle-necked by data. Template-based methods [87, 88, 7] cannot generalize to unseen objects as they are trained only for specific object templates. Template free methods like $\text{PC}^{2}$ [56] cannot separate human and object, and have limited shape accuracy. See Tab. 1 for detailed comparison. To alleviate these issues, we propose Hierarchical Diffusion Model (HDM), that predicts accurate object shapes and reasons about human-object contacts without using template meshes. Our key idea is to decompose the combinatorial interaction space into separate human and object sub-spaces while preserving the interaction context. We first use a diffusion model to predict human and object points jointly along with segmentation labels and then use two separate diffusion models with cross attention that further refine the shapes of human and objects.
@@
  • We introduce the first method for procedural interaction generation that can synthesize large amounts of interaction with diverse object shapes. With this, we generate a million interaction images with 21k different objects paired with clean 3D ground truth.
@@
  • Our dataset, code for data generation and reconstruction will be publicly released to foster future works.
@@ 2 Related Work
Interaction Capture. Capturing 3D interactions from monocular inputs is an emerging research field in recent years, with works that reconstruct hand-object interaction from RGB [39, 17, 28, 22, 92] or RGBD [10, 11, 26] input, or predict contacts from RGB images [78, 14, 32] and works that model human-scene interaction from single image [27, 43, 70, 8] or video [94, 24]. A recent line of works model full body interacting with dynamic large objects [60, 101, 25, 91, 49, 83, 37]. BEHAVE [7] captured the first dataset for benchmarking methods in this field, followed by more datasets [33, 97], which allow methods to reconstruct 3D human and object from single RGB images [98, 87, 82] or videos [88]. Despite impressive results, they require predefined mesh templates, which limits applicability to novel objects. In contrast, our method does not require templates and generalizes well to unseen objects.
Synthetic Datasets are powerful resources to pre-train large backbone networks. For 3D humans, synthetic scans [1, 4, 77, 59] are used extensively to train human reconstruction methods [66, 67, 89, 75, 18, 5, 6, 90]. Recent work BEDLAM [9] showed that training purely on synthetic datasets [9, 59] allows strong generalization. Orthogonal to these, large scale 3D object CAD model datasets [12, 86] are also used to pretrain backbone models[36, 95, 58, 47]. Other works [63, 23, 85, 20] consider generating diverse and photo-realistic scenes. While being useful in modelling humans, objects or scenes respectively, they do not consider interactions. Our proposed data generation approach can generate millions of interactions with diverse object shapes, allowing for training interaction reconstruction models with great generalization ability.
@@ 3.1.4 Dataset Rendering
Method Scalability.
We emphasize that the proposed procedural generation is a scalable solution that can generate large-scale datasets with only a small amount of effort for interaction: with 2k different interactions (e.g. BEHAVE [7] chair interaction), 6k different objects (e.g. Shapenet chairs [12]) and 100 human scans (e.g. MGN [4]), one can have maximum 1.2 billion different variations in total, which is not possible with real data capture. The data scale allows for training powerful models that reach performance not obtainable by training on real data only. An example of such a method is detailed in the next section.
@@ 3.2.1 Preliminaries
Task Overview. Given an input RGB image $\mathbf{I}$ of a person interacting with an object, we aim to jointly reconstruct 3D human and object point clouds $\mathbf{P}^{h},\mathbf{P}^{o}$. Same as prior works [87, 98, 88], we assume known 2D human and object segmentation masks, which we consider a weak assumption, given recent advances in 2D segmentation [40, 41, 71].
@@ 3.2.3 Hierarchical Diffusion for Interaction
Model Combination. With the separate networks $p_{\theta}^{h},p_{\theta}^{o}$, one can run the full reverse diffusion process from $t=T$ to $t=0$ and then combine the denoised points. However, this does not leverage the predicted interaction context from the joint reconstruction stage and is slow. We hence start the reverse diffusion steps from an intermediate step $t=T_{0}$ instead of $T$. Specifically, after denoising and segmentation with the joint model, we apply the forward diffusion process to $\mathbf{P}^{h}$ and $\mathbf{P}^{o}$ until step $t=T_{0}$ using Eq. 3. Then, the individual diffusion models take the noised points as input and gradually denoise them until step $t=0$. The forward process destroys local noisy predictions but keeps the global structure of human-object interaction. We empirically set $T_{0}=\frac{T}{2}$, see supp. for analysis of this value. We show in Tab. 5 and Fig. 5 that our hierarchical design is important to obtain sharp predictions.
@@ 4 Experiments
Data generation. We leverage the BEHAVE [7], InterCap [33], ShapeNet [12], Objaverse [19], ABO [15] and MGN [4] dataset to generate our synthetic data ProciGen. BEHAVE and InterCap capture multi-view images of humans interacting with 20 and 10 different objects respectively. ShapeNet [12] and Objaverse [19] are large-scale datasets that provide 3D object models as meshes with textures. ABO [15] is a smaller shape dataset with high-quality material and textures. The objects from ShapeNet and ABO are aligned in canonical space while objects from Objaverse are not aligned. MGN [4] is a dataset consisting of 100 human scans paired with SMPL-D registration that allows reposing scans while preserving clothing deformation.
Following the same train-test split from [88], we randomly sample from 380k interactions in BEHAVE and InterCap training set, 21k different shapes in ShapeNet, ABO and Objaverse, and 100 different human shapes and textures in MGN. In total, we generate $\sim$1.1million training images. Please see supplementary for more data distribution details.
@@ 4.1 Reconstruction on BEHAVE and InterCap
We compare our method with CHORE [87] and $\text{PC}^{2}$[56] on BEHAVE[7] and InterCap [33] test set in Tab. 2 and Fig. 4. We train CHORE and $\text{PC}^{2}$ on the training set of BEHAVE and InterCap. Our HDM is trained on our synthetic ProciGen with or without BEHAVE and InterCap training set.
@@ 4.2 Contribution of our ProciGen and HDM
We propose ProciGen for interaction data generation and HDM for interaction reconstruction. To decouple the contribution of our data and method, we compare our method against $\text{PC}^{2}$ [56] trained on BEHAVE [7] only (Tab. 3 a-b) and BEHAVE + our ProciGen dataset (Tab. 3 c-d). The methods are evaluated on the BEHAVE test set. It can be seen that both our proposed data and model are important to obtain the most accurate reconstruction.
@@ 4.3 Generalization Performance
Some qualitative results are shown in Fig. 5. Our method reconstructs human and object separately with high shape fidelity. We also show the generalization results to COCO dataset [45] in Fig. 6. Our method trained only on our ProciGen data generalizes well to in-the-wild images with large object shape variations. See Supp. for more examples.
@@ 4.4 Ablating the Hierarchical Diffusion Model
Our HDM predicts interaction semantics and better shape fidelity. In Table 5, we ablate other alternatives to our method on the chair category from BEHAVE test set [7] (824 images) due to resource limit. All methods are trained on our ProciGen dataset.
The human-object segmentation allows us to compute the contacts and manipulate human and object separately. An alternative is projecting the predicted points to 2D image and segment points based on the masks. Due to occlusion and complex interaction, this segmentation is inaccurate, which is reflected in the large separate human and object errors in Tab. 5 a. The model that predicts human, object and segmentation with a single model (Tab. 5 b) also does not work as it is difficult to learn high-fidelity interaction shapes within one model.
@@ 5 Conclusion
We train our method with the proposed synthetic dataset and evaluate it on BEHAVE and InterCap datasets. Results show that our method significantly outperforms CHORE which requires template meshes and $\text{PC}^{2}$ which does not reason interaction semantics. Ablation studies also show that our synthetic dataset is important to boost the performance and generalization ability of both $\text{PC}^{2}$ and our model. Our method generalizes well to real images from COCO that have diverse object geometries, which is a promising step toward real in-the-wild reconstruction. Our code will be released to promote future research.
Acknowledgements. We thank RVH group members [2], especially Yuxuan Xue, for their helpful discussions. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 409792180 (Emmy Noether Programme, project: Real Virtual Humans), and German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. Gerard Pons-Moll is a Professor at the University of Tübingen endowed by the Carl Zeiss Foundation, at the Department of Computer Science and a member of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645.
@@ A.1 ProciGen Data Generation
Figure 8: Distribution of interactions per object category. Our dataset features interaction data with very diverse object shapes, which is not possible via real data capture
@@ A.2 HDM: Hierarachical Diffusion Model
The first equation projects origin to $\mathbf{p}_{c}$ and the second equation projects $\mathbf{p}_{e}$ to the middle right of the 2D crop $\mathbf{p}^{\text{2D}}_{e}=(c_{x}+w/2,c_{y})$. This is a linear system of four equations with four unknowns (camera translation and depth $z$), leading to a unique solution for the translation $\mathbf{t}_{c}$. We empirically set $\sigma$ to different values for different categories based on the estimation error on the BEHAVE training set. From Fig. 6, Fig. 15, Fig. 14, Fig. 16, Fig. 17 and , Fig. 18, it can be seen that our method can reconstruct human and object well on different datasets using our estimated translation.

### 06 Apr 2024 v3

@@ Abstract
Reconstructing human-object interaction in 3D from a single RGB image is a challenging task and existing data driven methods do not generalize beyond the objects present in the carefully curated 3D interaction datasets. Capturing large-scale real data to learn strong interaction and 3D shape priors is very expensive due to the combinatorial nature of human-object interactions. In this paper, we propose ProciGen (Procedural interaction Generation), a method to procedurally generate datasets with both, plausible interaction and diverse object variation. We generate 1M+ human-object interaction pairs in 3D and leverage this large-scale data to train our HDM (Hierarchical Diffusion Model), a novel method to reconstruct interacting human and unseen object instances, without any templates. Our HDM is an image-conditioned diffusion model that learns both realistic interaction and highly accurate human and object shapes. Experiments show that our HDM trained with ProciGen significantly outperforms prior methods that require template meshes, and our dataset allows training methods with strong generalization ability to unseen object instances. Our code and data are released.
@@ 1 Introduction
Modelling interactions between humans and their surroundings is important for applications like creating realistic avatars, robotic control and gaming. In this paper, we address the task of jointly reconstructing human and object from a monocular RGB image, without any prior object templates. This is very challenging due to depth-scale ambiguity, occlusions, diverse human pose and object shape variations. Data-driven methods have shown great progress in reconstructing humans [69, 70, 71, 46, 41, 31, 85] or objects [55, 112] from monocular inputs thanks to large-scale datasets [9, 37, 63, 1, 108, 19, 12, 94]. However, methods for joint interaction reconstruction are still constrained by the amount of available data. Recent datasets like BEHAVE [7], InterCap [35] capture real interactions with 10 to 20 different objects, which is far away from the number of objects in reality: the chair category from ShapeNet [12] alone has more than 6k different shapes. Training on these real datasets has limited generalization ability to unseen objects (Sec. 4.3). Capturing real interaction data with more objects is prohibitively expensive due to the combinatorial nature: the number of humans times the number of objects leads to a huge number of variations. This motivates us to generate synthetic data which has been shown effective for pre-training reconstruction methods [9, 63, 29, 70, 55].
@@
Synthesizing realistic interaction for different objects is non-trivial due to variations of object topology, geometry details and complex interaction patterns. To address this, we propose ProciGen: Procedural interaction Generation, a method to generate interaction data with diverse object shapes. We design our method based on the key idea that the way humans interact with objects of the same category is similar. And despite the geometry variations, one can still establish semantically meaningful correspondence between different objects. More specifically, we train an autoencoder to obtain correspondences between different objects from the same category, which are then used to transfer contacts from already captured human-object interactions to new object instances. Our method is scalable and allows the multiplicative combination of datasets to generate over a million interactions with more than 21k different object instances, which is not possible via real data capture.
@@
Current reconstruction methods [95, 96, 7, 60] are not only bottle-necked by data. Template-based methods [95, 96, 7] cannot generalize to unseen objects as they are trained only for specific object templates. Template free methods like $\text{PC}^{2}$ [60] cannot separate human and object, and have limited shape accuracy. See Tab. 1 for detailed comparison. To alleviate these issues, we propose HDM: Hierarchical Diffusion Model, that predicts accurate shapes and reasons about human-object semantics without using templates. Our key idea is to decompose the combinatorial interaction space into separate human and object sub-spaces while preserving the interaction context. We first use a diffusion model to jointly predict human, object and segmentation labels, and then use two separate diffusion models with cross attention that further refine the separate predictions.
@@
  • We introduce the first procedural interaction generation method for synthesizing large-scale interaction data with diverse objects. With this, we generate 1M+ interaction images with 21k+ objects paired with 3D ground truth.
@@
  • Our dataset and code are publicly released.
@@ 2 Related Work
Interaction Capture. Modelling 3D interactions has been an emerging research field in recent years, with works that model hand-object interaction from RGB [42, 17, 29, 22, 102], RGBD [10, 11, 26] or 3D [114, 116, 80, 42, 83] input, or predict contacts from RGB images [86, 14, 34] and works that model human-scene interaction from single image [28, 47, 74, 8, 101] or video [105, 24]. A recent line of works model full body interacting with dynamic large objects [64, 113, 25, 100, 53, 91, 40, 39, 78, 27, 44]. BEHAVE [7] and follow up works [35, 109] capture interaction datasets, which allow training and benchmarking methods [97] for reconstructing 3D human-object from single RGB images [110, 95, 90] or videos [96]. Despite impressive results, they require predefined mesh templates, which limits applicability to new objects. Our method is template-free and generalizes well to unseen objects.
Synthetic Datasets are powerful resources to deep networks. For humans, synthetic rendering of 3D scans [1, 4, 84, 63, 103] are used extensively to train human reconstruction methods [70, 71, 98, 81, 18, 5, 6, 99]. Recent work BEDLAM [9] showed that training purely on synthetic datasets [9, 63] allows strong generalization. Orthogonal to these, large scale 3D object CAD model datasets [12, 94] are also used to pretrain backbone models[38, 107, 62, 51]. Other works [67, 23, 93, 20] consider generating diverse scenes. While being useful for humans, objects or scenes respectively, they do not consider interactions. Our proposed approach can generate millions of interactions with diverse object shapes, allowing for training interaction reconstruction models with great generalization ability.
@@ 3.1.4 Dataset Rendering
Method Scalability. We emphasize that the proposed procedural generation is a scalable solution that can generate large-scale datasets with only a small amount of effort for data capture: with 2k different interactions (e.g. BEHAVE [7] chair interaction), 6k different objects (e.g. Shapenet chairs [12]) and 100 human scans (e.g. MGN [4]), one can have maximum 1.2 billion different variations in total, which is not possible with real data capture. The data scale allows for training powerful models that reach performance not obtainable by training on real data only. An example of such a method is detailed in the next section.
@@ 3.2.1 Preliminaries
Task Overview. Given an input RGB image $\mathbf{I}$ of a person interacting with an object, we aim to jointly reconstruct 3D human and object point clouds $\mathbf{P}^{h},\mathbf{P}^{o}$. Same as prior works [95, 110, 96], we assume known 2D human and object masks, which we consider a weak assumption, given recent advances in 2D segmentation [43, 45, 75].
@@ 3.2.3 Hierarchical Diffusion for Interaction
Model Combination. With the separate networks $p_{\theta}^{h},p_{\theta}^{o}$, one can run the full reverse diffusion process from $t=T$ to $t=0$ and then combine the denoised points. However, this does not leverage the predicted interaction context from the joint reconstruction stage and is slow. We hence start the reverse diffusion steps from an intermediate step $t=T_{0}$ instead of $T$. Specifically, after denoising and segmentation with the joint model, we apply the forward diffusion process to $\mathbf{P}^{h}$ and $\mathbf{P}^{o}$ until step $t=T_{0}$ using Eq. 3. Then, the individual diffusion models take the noised points as input and gradually denoise them until step $t=0$. The forward process destroys local noisy predictions but keeps the global structure of human-object interaction. We set $T_{0}=\frac{T}{2}$, see supp. for analysis of this value. Our hierarchical design is important to obtain sharp predictions, see Tab. 5 and Fig. 5.
@@ 4 Experiments
Data generation. We leverage the BEHAVE [7], InterCap [35], ShapeNet [12], Objaverse [19], ABO [15] and MGN [4] dataset to generate our synthetic data ProciGen. BEHAVE and InterCap capture multi-view images of humans interacting with 20 and 10 different objects respectively. ShapeNet [12], Objaverse [19] and ABO [15] provide 3D object models as meshes paired with textures. The objects from ShapeNet and ABO are aligned in canonical space while objects from Objaverse are not aligned. MGN [4] consists of 100 human scans paired with SMPL-D registration that allows reposing scans while preserving clothing deformation.
Following the same split from [96], we randomly sample from 380k interactions in BEHAVE and InterCap training set, 21k different shapes in ShapeNet, ABO and Objaverse, and 100 different human shapes and textures in MGN. In total, we generate $\sim$1.1million training images. Please see supplementary for more data distribution details.
@@ 4.1 Reconstruction on BEHAVE and InterCap
We compare our method with CHORE [95] and $\text{PC}^{2}$[60] on BEHAVE[7] and InterCap [35] test set in Tab. 2 and Fig. 4. We train CHORE and $\text{PC}^{2}$ on the training set of BEHAVE and InterCap. Our HDM is trained on our synthetic ProciGen with or without BEHAVE and InterCap training set. We also report per-category accuracy in supplementary.
@@ 4.2 Contribution of our ProciGen and HDM
We propose ProciGen for interaction data generation and HDM for interaction reconstruction. To decouple the contribution of our data and method, we compare our method against $\text{PC}^{2}$ [60] trained on BEHAVE [7] only (Tab. 3 a-b) and BEHAVE + our ProciGen dataset (Tab. 3 c-d). The methods are evaluated on the BEHAVE test set. It can be seen that both our proposed data and model are important to obtain the most accurate reconstruction. We also report the model performance vs. data amount in supplementary.
@@ 4.3 Generalization Performance
Some qualitative results are shown in Fig. 5. Our method reconstructs human and object with high shape fidelity. We also show the generalization results to COCO dataset [49] in Fig. 6. Our method trained only on our ProciGen data generalizes well to in-the-wild images with large object shape variations. See Supp. for more generalization examples.
@@ 4.4 Ablating the Hierarchical Diffusion Model
Our HDM predicts interaction semantics and better shapes. In Tab. 5, we ablate alternative designs to our method on the 824 chair images from BEHAVE test set [7] due to resource limit. All methods are trained on our ProciGen dataset.
The human-object segmentation allows us to compute the contacts and manipulate human and object separately. An alternative is projecting the predicted points to 2D image and segment points based on the masks. Due to occlusion and complex interaction, this segmentation is inaccurate, as reflected in the large human and object errors in Tab. 5a. The model that predicts human, object and segmentation with a single model (Tab. 5 b) also does not work as it is difficult to learn high-fidelity interaction shapes.
@@ 5 Conclusion
We train our method with the proposed synthetic dataset and evaluate it on BEHAVE and InterCap datasets. Results show that our method significantly outperforms CHORE which requires template meshes and $\text{PC}^{2}$ which does not reason interaction semantics. Ablation studies also show that our synthetic dataset is important to boost the performance and generalization ability of both $\text{PC}^{2}$ and our model. Our method generalizes well to real images from COCO that have diverse object geometries, which is a promising step toward real in-the-wild reconstruction. Our code and data are released to promote future research.
Acknowledgements. We thank RVH group members [2], especially Yuxuan Xue, for their helpful discussions. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 409792180 (Emmy Noether Programme, project: Real Virtual Humans), and German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A, and Amazon-MPI science hub. Gerard Pons-Moll is a Professor at the University of Tübingen endowed by the Carl Zeiss Foundation, at the Department of Computer Science and a member of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645.
@@ A.1 ProciGen Data Generation
Figure 8: Distribution of interactions per object category. Our dataset features interaction data with very diverse object shapes, which is not possible via real data capture.
@@ A.2 HDM: Hierarachical Diffusion Model
The first equation projects origin to $\mathbf{p}_{c}$ and the second equation projects $\mathbf{p}_{e}$ to the middle right of the 2D crop $\mathbf{p}^{\text{2D}}_{e}=(c_{x}+w/2,c_{y})$. This is a linear system of four equations with four unknowns (camera translation and depth $z$), leading to a unique solution for the translation $\mathbf{t}_{c}$. We empirically set $\sigma$ to different values for different categories based on the estimation error on the BEHAVE training set. Furthermore, we compute $\mathbf{p}_{c}$ as the centroid of all 2D points inside the human and object masks. From Fig. 6, Fig. 17, Fig. 16, Fig. 18, Fig. 19 and , Fig. 20, it can be seen that our method can reconstruct human and object well on different datasets using our estimated translation.
A.3 Dataset statistics
We generate our ProciGen dataset based on interactions from BEHAVE [7] and InterCap [35], human scans from MGN [4], object shapes from ShapeNet [12], Objaverse [19] and ABO [15]. When generating our data, we mainly consider the variation of object shapes and interaction poses while the object sizes remain the same. We also try to avoid large imbalances among objects. Therefore, chairs and tables are two dominant categories as they have the most geometry and interaction pose variations (Fig. 7, Fig. 8). Other categories have similar amounts of synthetic data as they have similar amounts of interaction poses. The difference comes from failures in joint optimization due to irregular mesh.
In total we rendered 1.1M interaction images with 21555 different object shapes. The distribution for object shapes and interactions per category are shown in Fig. 7 and Fig. 8. Our dataset has very diverse object shapes, especially for chairs and tables whose geometry also varies a lot in reality. Our procedural generation method is a scalable solution and it allows for generating large-scale interaction datasets with great amount of variations which is not obtainable via capturing real data.
Appendix B Additional Experiments and Results
B.1 Per-category reconstruction accuracy
We report the accuracy of each category in Tab. 6. Our method consistently outperforms baselines in almost all categories. While the improvements in numbers look small, the visual difference is quite significant, as shown in the paper Fig. 4, Fig. 6.
Method CHORE $\uparrow$ PC2 $\uparrow$ Ours $\uparrow$
Hum.
Obj.
Comb.
Comb.
Hum.
Obj.
Comb.
Chair
0.373
0.491
0.443
0.407
0.384
0.521
0.463
Ball
0.330
0.388
0.374
0.424
0.395
0.517
0.471
Backpack
0.399
0.509
0.469
0.436
0.397
0.457
0.444
Table
0.304
0.455
0.389
0.470
0.379
0.642
0.517
Basket
0.301
0.266
0.292
0.381
0.412
0.297
0.364
Box
0.352
0.347
0.362
0.409
0.414
0.401
0.424
Keyboard
0.335
0.412
0.383
0.450
0.353
0.606
0.493
Monitor
0.358
0.412
0.395
0.377
0.368
0.348
0.370
Suitcase
0.400
0.477
0.443
0.404
0.431
0.484
0.462
Stool
0.351
0.479
0.424
0.443
0.394
0.543
0.479
Toolbox
0.281
0.330
0.306
0.398
0.373
0.400
0.403
Trashbin
0.376
0.402
0.398
0.387
0.407
0.414
0.422
BEHAVE all
0.345
0.426
0.397
0.423
0.392
0.498
0.457
Chair
0.389
0.468
0.433
0.470
0.384
0.604
0.500
Cup
0.412
0.538
0.510
0.566
0.487
0.601
0.578
Skateboard
0.520
0.684
0.612
0.578
0.491
0.739
0.624
Bottle
0.426
0.501
0.495
0.592
0.549
0.582
0.593
InterCap all
0.406
0.513
0.469
0.506
0.440
0.607
0.534
Table 6: Per-category F-score@0.01m comparison. Note that PC2 cannot separate human-object hence we only report the combined error, and that CHORE requires template meshes. Our method outperforms baselines for almost all categories.
B.2 Performance vs. amount of data
We show in Table 3 that our data contributes a lot to improve the reconstruction accuracy. To further understand the data contribution, we train our model for the same epochs with different amounts of synthetic data and test on BEHAVE images without fine-tuning. The performance vs. data plot is shown in Fig. 9. More data consistently leads to better performance both quantitatively and qualitatively.
Figure 9: Reconstruction performance vs. amount of data. It can be seen that more data leads to better results.
B.3 Analysis of $T_{0}$ for our HDM